# Backdoor-based Explainable AI Benchmark for High Fidelity Evaluation of Attribution Methods

### v1

@@ 5.3 Output Choice
In general, contemporary attribution methods lack a clear distinction between using model logits and softmax probabilities as the model output for computing the attributions [27]. This can compromise benchmarking transparency. In Fig. 5, we compare the attack success rate and trigger recall on BackX benchmark when attributions use output logits and probabilities on CIFAR-10, GTSRB, and ImageNet. Following the conventions from Fig. 4, we report the differences between the values.
The results demonstrate that CAM-based methods tend to rely on explaining probabilities to achieve better performance, whereas integration-based methods prefer explaining output logits. Gradient-based methods exhibit a significant fluctuation between explaining logits and probabilities. Moreover, explaining logits leads to stronger localization capabilities in both gradient-based and integration-based attribution methods, as implied by the results in Fig. 5(b). This is indicative of the inherent richness of raw logits in facilitating class localization.
It is worth noting that the attribution differences between the choices of logits and probabilities become narrower as the number of classes increases. This convergence is attributed to softmax normalization, which redistributes probability mass more evenly across categories. This phenomenon is mathematically evident in the softmax function, where the denominator $\sum_{j=1}^{N}e^{z_{j}}$ increases with the total number of categories $N$, leading to the convergence of probabilities.
[…]
@@
Observation 2 CAM-based methods rely on output probabilities to gather distinctive class information to construct accurate activation maps. In contrast to output probabilities, explaining logits enables attributions to preserve unnormalized original activations for a class, leading to enhanced localization capability. However, the disparity of attributions between explaining logits and probabilities ultimately diminishes as the number of classes increases.
@@
Guided by this observation, we use CAM-based methods to explain output probabilities, while we use both gradient-based and integration-based attribution methods to explain logits in our subsequent experiments.
5.4 Re-Calibrating and Contrastive Output
Recent research has explored alternatives to enhance reliability of baseline attribution results [27], [28]. Here, we conduct comparative experiments for such choices. Re-calibrated attributions [27] have been introduced to enhance integration-based methods by estimating attributions from identified valid references. In Fig. 7, we compare the re-calibrated attributions and original attributions across integration-based attribution methods, including IG-SG, IG-Uni, AGI and LPI. It is observed that re-calibrated attributions further enhance the performance of the original attributions on CIFAR-10, GTSRB and ImageNet datasets. In addition, the performance of re-calibrated attributions is observed to degrade with decreasing mean input pixel values, aligning with our Observation 2. This demonstrates the sensitivity of integration-based methods to variations in the reference values. Despite the enhancement achieved through re-calibration, it remains unique to integration-based attribution methods. To maintain evaluation consistency across different methods, we retain the use of original attributions in the integration methods for the subsequent experiments.
Figure 8 offers a comparison between attributions computed by our established consistent setup and contrastive outputs [27] on CIFAR-10, GTSRB, and ImageNet. It is notable that explaining contrastive output yields similar performance compared to explaining manually selected outputs, particularly evident in cases involving FullGrad and integration-based methods. This observation underscores the superiority of contrastive output over mutual selection. However, explaining contrastive output in CAM and Gradient-based methods does not achieve comparable performance. The following observation is derived from the above analysis.
Observation 3 The re-calibration technique consistently improves the performance of integration-based attribution methods. Attribution derived from explaining contrastive model output can achieve competitive results, particularly for CAM-based and integration-based methods.
@@ 5.4 Re-Calibrating and Contrastive Output
In summary, our experimental results demonstrate that different setups yield distinctive performances for the same attribution method. This implies that the performance of these methods is susceptible to misinterpretation. We offer clear guidelines for consistent and fair benchmarking for the considered methods. To the best of our knowledge, there has been no prior research addressing the need for a standardized consistent setup. We advocate for the adoption of a unified configuration, which is imperative for the advancement of attribution methods.
@@
Figure 9: Benchmarking of attribution methods under Blend attack using the attack success rate metric. Lower curves indicate better results.
Figure 10: Benchmarking of attribution methods under Blend attack using the trigger recall metric. Higher curves indicate better results.
@@ 6 Benchmarking
In Fig. 9, we provide results for Trojaned ResNet-18 models using the Blend attack with trigger visibility of $0.5$. We employ nine attribution methods with varying image recovery rates across three datasets: ImageNet, CIFAR-10 and GTSRB. The corresponding trigger recall, as the image recovery rate increases, is reported in Fig. 10. Experimental results reveal that integration-based methods and gradient-based methods generally outperform CAM-based methods. Integration-based methods achieve superior performance on GTSRB, while gradient-based methods outperform others on CIFAR-10. Importantly, integration-based methods exhibit higher stability across different datasets.
@@
Figure 11: Benchmarking of attribution methods under ISSBA attack using the attack success rate metric. Lower curves indicate better results.
Figure 12: Comparison of fractional logit change between attribution methods under Blend attack. Bubble size is scaled by the method’s performance. Approaching top-left corner in each plot indicates superior performance.
Figure 13: Results of trigger recall changes with trigger visibility between attribution methods. Mean curves over various methods are highlighted. Higher trigger recall values correspond to an increased capacity to defend against backdoor attacks.
Figure 14: Comparison of fractional logit change between attribution methods under ISSBA attack. Bubble size is scaled by the method’s performance. Approaching top-left corner in each plot indicates superior performance.
@@
It is observed that GGCAM significantly enhances CAM across all datasets with element-wise estimated attributions by incorporating Grad. Moreover, attribution methods estimated from perturbed inputs with steep slope curves, such as SG and AGI, demonstrate the best performance in locating important features. However, AGI’s random class selection for calculating adversarial examples undermines its stability.
@@
Figure 11 shows the attack success rate on a Trojaned model through ISSBA attack [25] on CIFAR-10 and GTSRB. Due to input-specific triggers demonstrating weak attack capabilities on larger-scale input samples, as they can be easily recovered by random attributions, we have excluded ImageNet from our experiments. Fine-grained integration-based and gradient-based methods can achieve outstanding performance for detecting invisible triggers. CAM-based methods, which fail to capture element-wise triggers, still manage to achieve competitive results compared to the integration-based and gradient-based methods.
@@
In Fig. 12 and Fig. 14, we compare the fractional logit change for both the Blend and ISSBA attacks. The experimental results show the fractional change in logit output of both the true class $y$ and the target class $\tilde{y}$ to create a bubble chart. A faithful attribution method is expected to reduce the output of target class $\tilde{y}$ while recovering the output of true class $y$. The markers representing different attribution methods are scaled by the corresponding performance. In this metric, integration-based methods exhibit both high performance and consistency. Additionally, compared to recovering the output of the true class $y$, all attribution methods are effective at decreasing the output of the target class $\tilde{y}$. Notably, CAM-based attribution methods can achieve competitive results in recovering the output of $y$ compared to $\tilde{y}$, revealing the weak attack capability of invisible attacks. The results indicate that the attribution methods tend to perturb the target class rather than fully recover it. More results on fractional probability change and attribution visualizations are provided in Appx. 13. We make the following observation based on the overall results.
@@
Observation 4 Gradient-based attribution methods tend to achieve better results, while integration-based methods exhibit more stability across different datasets. As compared to recovering the original class $y$, all attribution methods are better at reducing the misclassification to target $\tilde{y}$.
@@ 7 Backdoor Defense with Attributions
Here, we investigate the potential of the analyzed attribution methods for defending against backdoor attacks. In Fig. 13, we assess the change in attribution performance, measured by trigger recall, across varying trigger visibility. Counter to general intuition, we observe that attribution methods do not lead to better trigger localization with higher visibility of the trigger. This unanticipated behavior can be attributed to the fact that learning a robust feature, such as a clear trigger, does not demand fine-grained tuning of the model weights to attract its attention. In other words, better trigger visibility actually leads to a relatively easier adversarial learning objective. Hence, trigger visibility fails to show a positive correlation with trigger recall through attribution. In Appx. 14.2, more visualization results of attribution maps are presented concerning the detection of triggers with varying degrees of visibility.
Figure 15: Results of trigger recall changes with poisoning rate. Higher values correspond to better defense capability.
Figure 15 also reports the change in trigger recall as the poisoning rate varies. Interestingly, the poisoning rate also has only a limited effect on the performance of attribution methods. We provide more experimental results on defending backdoor attacks with attribution methods in Appx. 14. Based on our experiments, we make the following observation to provide guidance for employing attribution in defense against backdoor attacks.
Observation 5 Fixed or input-specific invisible attacks do not necessarily render defense more challenging. Fine-grained perturbation-based attribution methods, while incurring a higher computational cost, are more useful for backdoor defense. Additionally, CAM-based methods guided by fine-grained attributions are better choices for defense.
@@ 8 Conclusion
To establish a reliable benchmark for eXplainable Artificial Intelligence (XAI) in attribution methods, we devised a set of fidelity criteria aimed at assessing the evaluation faithfulness of such benchmarks. This facilitated a thorough comparison of existing benchmarks. Subsequently, a Backdoor-based evaluation framework (BackX) is developed, which underwent theoretical validation to ensure a superior level of fidelity criteria. Leveraging this fidelity benchmark, our study systematically explored various attribution methods, revealing their distinctive properties. This systematic exploration allowed for a consistent setup across different attribution methods, thereby mitigating confounding factors inherent in attribution estimation. Our framework employs this setup to conduct comprehensive benchmarking of attribution methods, highlighting both their strengths and weaknesses.
Appendix
9 Notation
In Table II, we provide a summary of the notations used in the paper, along with their corresponding definitions.
TABLE II: The used notations and the corresponding definition.
@@ 9 Notation
$S^{(k)},S^{*}\in\{0,1\}^{n}$ Attribution mask indicating $k$most important elements and a mask of a trigger pattern
@@
$\boldsymbol{{\rm I}}(\cdot)$ Bool function
@@
$A^{k}_{i,j}$ ($i$-th,$j$-th) element of activations in $k$-th layer
@@
10 Proof
@@
Assuming that $\tilde{\epsilon}$ is a subset of $\epsilon$, any deviation from the original data distribution caused by the poisoned sample $\tilde{x}$ is constrained within the boundary of $\epsilon$. The recovery process aims to undo the perturbation introduced by the poisoned sample $\tilde{x}$. By constraining the transformation from $\tilde{x}$ to $\hat{x}$ such that $\Omega(x,\hat{x})\leq\Omega(x,\tilde{x})$, the recovered sample $\tilde{x}$ can be ensured to remain within the boundary defined by $\epsilon$. Therefore, under the condition that $\tilde{\epsilon}\leq\epsilon$, the recovery process preserves the original data distribution, and the recovered sample $\hat{x}$ is preserved within the original distribution. Thus, if $\tilde{\epsilon}\leq\epsilon$, the recovered sample $\hat{x}$ is preserved within the original data distribution.
@@
Given a recovery image $\hat{x}$, the performance of attribution methods for a target class $c$ evaluated in our BackX benchmark can be represented as $I(\hat{x};c)$. We follow the derivation on the multi-information as Vergara & Estévez [58], Rong et al. [17]. Assume an attribution mask $S$, the multi-information $I(\hat{x};c;M)$ can be represented as
@@
$I(\hat{x};c)=I(c;\hat{x}|S)+I(S;c)-I(c;S|\hat{x}),$ (13)
where $I(c;\hat{x}|S)$ represents the target we aim to estimate, representing the mutual information between the recovery input $\hat{x}$ and a class $c$ for a given mask $S$. $I(S;c)$ represents the mutual information between $S$ and $c$, which we aim to eliminate. The last term $I(c;S|\hat{x})$ indicates mutual information between $S$ and $c$ given $\hat{x}$, compensating for $I(S;c)$.
Minimizing the mutual information $I(S;c)$ between the mask and the class label leads to the approximation that $I(S;c)$ approaches $I(c;S|\hat{x})$, specifically $I(S;c)\approx I(c;S|\hat{x})$. Thus, this mitigation of the leaked information $I(S;c)$ from the mask $S$ leads to an improvement in the reliability of the evaluation results for $I(\hat{x};c)$, specifically $I(\hat{x};c)\approx I(c;\hat{x}|S)$. ∎

### v2

@@ 4.2 Output Choice
In general, contemporary attribution methods lack a clear distinction between using model logits and softmax probabilities as the model output for computing the attributions (Wang & Wang, 2022). This can compromise benchmarking transparency. In Fig. 4, we compare the ASR and TR on BackX benchmark when attributions use output logits and probabilities on CIFAR-10, GTSRB, and ImageNet. Following the conventions from Fig. 3, we report the differences between the values. The results demonstrate that CAM-based methods tend to rely on explaining probabilities to achieve better performance, whereas integration-based methods prefer explaining output logits. Gradient-based methods exhibit a significant fluctuation between explaining logits and probabilities. In addition, explaining logits leads to stronger localization capabilities in both gradient-based and integration-based attribution methods, as shown in Fig. 4(b). However, it is worth noting that the attribution differences between the choices of logits and probabilities become narrower as the number of classes increases. We make the following observation from the experiments.
@@
Observation 2. CAM-based methods rely on output probabilities to gather distinctive class information to construct accurate activation maps. In contrast to output probabilities, explaining logits enables attributions to preserve unnormalized original activations for a class, leading to enhanced localization capability. However, the disparity of attributions between explaining logits and probabilities ultimately diminishes as the number of classes increases.
@@
Guided by this, we use CAM-based methods to explain output probabilities, while we use gradient-based and integration-based attribution methods to explain logits in our subsequent experiments.
@@
In App. A.8, we further investigate alternative contrastive outputs(Wang & Wang, 2022) and recalibrated attributions (Yang et al., 2023b), offering additional insights. In summary, our experimental results show that different setups yield distinctive performances for the same attribution method. This implies that the performance of these methods is susceptible to misinterpretation. We offer a standardized setup for consistent and fair benchmarking. To the best of our knowledge, there has been no prior research addressing the need for a standardized consistent setup. We advocate for the adoption of a unified configuration, which is imperative for the advancement of attribution methods.
@@
Figure 5: Benchmarking of attribution methods using the attack success rate metric. Lower is better.
Figure 6: Comparison of fractional logit change using (a) Blend attack and (b) ISSBA attack. Bubble size is scaled by overall performance. Bubbles near the top-left indicate better results.
@@ 5 Benchmarking
Figure 7: Benchmarking of attribution methods under Blend attack using trigger recall metric. Higher is better.
@@
In Fig. 5(a), we provide results for Trojaned ResNet-18 models using the Blend attack with trigger visibility of $0.5$. We employ nine attribution methods with varying image recovery rates across three datasets: ImageNet, CIFAR-10 and GTSRB. The corresponding trigger recall, as the image recovery rate increases, is reported in Fig. 7. Experimental results reveal that integration-based methods and gradient-based methods generally outperform CAM-based methods. Integration-based methods achieve superior performance on GTSRB, while gradient-based methods outperform others on CIFAR-10. Importantly, integration-based methods exhibit higher stability across different datasets. It is observed that GGCAM significantly enhances CAM across all datasets with element-wise estimated attributions by incorporating Grad. Moreover, attribution methods estimated from perturbed inputs with steep slope curves, such as SG and AGI, demonstrate the best performance in locating important features. However, AGI’s random class selection for calculating adversarial examples undermines its stability.
@@
Figure 5(b) shows the attack success rate on a Trojaned model through ISSBA attack (Li et al., 2021) on CIFAR-10 and GTSRB. Due to input-specific triggers demonstrating weak attack capabilities on larger-scale input samples, as they can be easily recovered by random attributions, we have excluded ImageNet from our experiments. Fine-grained integration-based and gradient-based methods can achieve outstanding performance for detecting invisible triggers. CAM-based methods, which fail to capture element-wise triggers, still manage to achieve competitive results compared to the integration-based and gradient-based methods.
@@
In Fig. 6, we compare the fractional logit change for both the Blend and ISSBA attacks. The experimental results show the fractional change in logit output of both the true class $y$ and the target class $\tilde{y}$ to create a bubble chart. A faithful attribution method is expected to reduce the output of target class $\tilde{y}$ while recovering the output of true class $y$. The markers representing different attribution methods are scaled by the corresponding performance. In this metric, integration-based methods exhibit both high performance and consistency. Additionally, compared to recovering the output of the true class $y$, all attribution methods are effective at decreasing the output of the target class $\tilde{y}$. Notably, CAM-based attribution methods can achieve competitive results in recovering the output of $y$ compared to $\tilde{y}$. The results indicate that the attribution methods tend to perturb the target class rather than fully recover it. More results on fractional probability change and attribution visualizations are provided in App. A.10. We make the following observation based on the overall results.
@@
Observation 3. Gradient-based attribution methods tend to achieve better results, while integration-based methods exhibit more stability across different datasets. As compared to recovering the original class $y$, all attribution methods are better at reducing the misclassification to target $\tilde{y}$.
More results on model architectures and model depths are provided in App. A.10.3. Beyond the vision domain, we extend BackX to language models, demonstrating its generality across modalities. Comprehensive benchmarking results on textual attribution are presented in App. A.11.
Figure 8: Results of trigger recall changes with (a) trigger visibility and (b) poisoning rate. Mean curves over various methods are highlighted. Higher trigger recall values correspond to an increased capacity to defend against backdoor attacks.
@@ 6 Backdoor Defense with Attributions
Here, we investigate the potential of the analyzed attribution methods for defending against backdoor attacks. In Fig. 8(a), we assess the change in attribution performance, measured by trigger recall, across varying trigger visibility. Counter to general intuition, we observe that attribution methods do not lead to better trigger localization with higher trigger visibility. This unanticipated behavior can be attributed to the fact that learning a robust feature, such as a clear trigger, does not demand fine-grained tuning of the model weights to attract its attention. In other words, better trigger visibility actually leads to a relatively easier adversarial learning objective. Hence, trigger visibility fails to show a positive correlation with trigger recall through attribution. Figure 8(b) also reports the change in trigger recall as the poisoning rate varies. Interestingly, the poisoning rate also has only a limited effect on the performance of attribution methods. Based on our experiments, we make the following observation to provide guidance of defending against backdoor attacks by attributions.
Observation 4. Counterintuitively, higher trigger visibility does not render improved defensibility, as highly visible triggers can steer predictions without requiring deep integration into the model’s reasoning process. Attribution methods maintain stable performance across poisoning rates, showing consistent feature learning modes for the same trigger pattern under varying attack scales.
Additional results on the resilience of attribution methods against various neural Trojans are in App. A.12, offering further insights into the effectiveness of attributions under diverse attacks.
@@ 7 Conclusion
To establish a reliable benchmark for XAI in attributions, we devised a set of fidelity criteria for evaluating attribution benchmarks, facilitating a thorough comparison of existing benchmarks. Subsequently, a backdoor-based evaluation framework (BackX) is developed, which underwent theoretical validation to ensure a superior level of fidelity criteria. Leveraging this fidelity benchmark, our study systematically explored various attribution methods, revealing their distinctive properties. This systematic exploration enabled a standardized setup for attribution estimation, mitigating confounding factors in benchmarking. BackX employs this setup to conduct comprehensive benchmarking of attributions, highlighting both their strengths and weaknesses. Furthermore, we provided insights into the defensibility of attributions against neural Trojans under varying attack settings.
Ethics Statement
This work contributes to the reliability and safety of AI systems by proposing a benchmark that rigorously evaluates the fidelity of attribution methods using controlled backdoor injections. Reliable attribution is essential for building trust, supporting accountability, and enabling safe deployment in sensitive domains such as healthcare and finance. By providing a principled way to reject unfaithful explanations, BackX helps prevent misleading interpretations of model behavior. While the use of backdoor techniques could raise dual-use concerns, our framework is strictly intended for controlled evaluation and does not facilitate malicious use. We believe this work offers a positive step toward more trustworthy and interpretable machine learning systems.
Reproducibility Statement
We provide the complete experimental setup in App. A.7, including trigger patterns, model configurations, hyperparameters, and software/platform details. Extended analyses and ablations are in Apps. A.8A.12. Formal proofs are provided in App. A.3. These sections contain the details needed to reproduce our results.
Appendix A Appendix
Contents of Appendix
A.1 Notation
In Table 2, we provide a summary of the notations used in the paper, along with their corresponding definitions.
Table 2: The used notations and the corresponding definition.
@@ A.1 Notation
$S^{(k)}$ Attribution mask indicating kk% most important elements
$S^{*}\in\{0,1\}^{n}$ Attribution binary mask of a trigger pattern
@@
$\bm{{\rm I}}(\cdot)$ Bool function
@@
$A^{k}_{i,j}$ (ii-th,$j$-th) element of activations in $k$-th layer
@@
A.2 Related Work
To explain model predictions, attribution methods assign importance scores to input features. Deconvnet (Zeiler & Fergus, 2014) and GBP (Springenberg et al., 2015) utilize deconvolution technique to compute feature importance. CAM (Zhou et al., 2016) and GradCAM (Selvaraju et al., 2017) calculate class activation maps for locating class-specific input features. Compared to CAM-based methods, InputGrad (Simonyan et al., 2014) calculates gradients with respect to the input for model explanation. SmoothGrad (Smilkov et al., 2017) aggregates input gradients across input samples with Gaussian noise, leading to a notable enhancement in localization ability. FullGrad (Srinivas & Fleuret, 2019) further integrates gradients from model biases, satisfying additional axioms. To guarantee more axioms,  Sundararajan et al. (2017) accumulate gradients from a reference to the input. To improve reference reliability, IG-SG (Smilkov et al., 2017), IG-SQ (Hooker et al., 2019) and IG-Uniform (Sturmfels et al., 2020) employ perturbed inputs as references. EG (Erion et al., 2021) employ training samples as references to maintain the distribution invariance. Pan et al. (2021) calculated class-specific adversarial samples as the reference. Yang et al. (2023b) recalibrated attributions with valid references. Wang & Wang (2022) explained a contrastive output for distinctive attributions, leading to class-contrastive attributions. Instead of explaining the model’s output,
Numerous explanation methods also have given rise to a multitude of XAI benchmarks. LeRF and MoRF (Samek et al., 2016) are extended from pixel flipping (Bach et al., 2015), which perturbs input samples to test attributions. Similarly, insertion and deletion games (Petsiuk et al., 2018) offer a benchmark for evaluating attributions of a single input sample. ROAR (Hooker et al., 2019) and DiffROAR (Shah et al., 2021) were proposed to retrain models on the perturbed images, ensuring models are learned within the distribution. ROAD (Rong et al., 2022) and DiffID (Yang et al., 2023a) offer ways to mitigate input distribution shift without costly model retraining. On the other hand, sensitivity-n (Ancona et al., 2018), SENSMAX and INFD (Yeh et al., 2019) focus on testing the fidelity of attributions by perturbing the input. Adebayo et al. (2018) randomized the model parameters and training labels to test the sanity of attributions. Other efforts have also been made to provide ground truth for the estimated attributions. DiFull, DiPart (Rao et al., 2022) and Pointing Game (Zhang et al., 2018) employ training annotations for evaluation. Khakzar et al. (2022) used model-optimized features for attribution ground truth. Lin et al. (2021) benchmarked explaining tools through analyzing the detection of visible triggers in input for Trojanned models. Additionally, Arras et al. (2022) employed the controlled VQA framework to generate synthetic images for benchmarking.
A.3 Proof
@@
Let $f$ be a benign model that correctly classifies input $x$ as label $y$, and maintains this prediction for all inputs $\bar{x}$ within an $\epsilon$-ball centered at $x$ under the distance metric $\Omega$. Formally, this assumption is expressed as
$\forall\bar{x}\in\mathbb{R}^{n},\,\Omega(x,\bar{x})\leq\epsilon\implies f(\bar{x})=f(x),$ (4)
where $\Omega(x,\bar{x})$ quantifies the distance between $x$ and $\bar{x}$.
[…]
$\Omega(x,\hat{x})\leq\tilde{\epsilon}\leq\epsilon,$ (5)
which implies that $\hat{x}$ lies within the $\epsilon$-ball centered at $x$. By the assumption in Eq. equation 4, it then holds that $f(\hat{x})=f(x)$.
Therefore, if the reconstruction error is bounded by the perturbation magnitude $\tilde{\epsilon}$ and $\tilde{\epsilon}\leq\epsilon$, the reconstruction process yields a sample $\hat{x}$ that remains within the model’s trusted prediction region. In this case, $\hat{x}$ preserves the original classification, effectively restoring the model’s behavior on Trojaned inputs. ∎
@@
Given a recovery image $\hat{x}$, the performance of attribution methods for a target class $c$ evaluated in our BackX benchmark can be represented as $I(\hat{x};c)$. We follow the derivation on the multi-information as Vergara & Estévez (Vergara & Estévez, 2014), Rong et al. (Rong et al., 2022). Assume an attribution mask $S$, the multi-information $I(\hat{x};c;M)$ can be represented as
@@
[…]
A.4 Benchmark Fidelity Discussion
In Tab. 1, we present a comparative analysis of different benchmarks with regard to their fidelity criteria assurance. Below, we discuss the details of the ratings provided in the table following three categories of the methods based on their underlying techniques.
Input Perturbation-based Methods. Since attribution maps are expected to highlight important input features, attribution benchmarks, including MoRF, LeRF (Samek et al., 2016), Ins.&Del. Games (Petsiuk et al., 2018), DiffID (Yang et al., 2023a) and ROAD (Rong et al., 2022), assess attribution methods by iteratively replacing pixels with zero or noise pixels. However, the input change causes an input distribution shift (Hooker et al., 2019), compromising the reliability of evaluation outcomes. To ensure input distribution invariance, ROAR (Hooker et al., 2019) and DiffROAR (Shah et al., 2021) retrain the model using perturbed input samples to incorporate the perturbed input samples within the training distribution. However, retraining models violate the criterion of functional mapping invariance. Thus, input perturbation-based attribution benchmarks shown in columns 1-3 of Tab. 1 struggle to ensure invariance to input distribution and the explained model.
Sanity and Sensitivity Checks. Sanity and sensitivity checks are originally proposed to test the fidelity of attribution methods for the explained input samples and models. Benchmarks such as sensitivity-n (Ancona et al., 2018), SENSMAX, and INFD (Yeh et al., 2019) aim to assess the sensitivity of attribution methods under input perturbations. While these benchmarks provide valuable assessment targets, these targets are unachievable by attribution methods in practice, resulting only in a partial guarantee of attribution verifiability. Adebayo et al. (Adebayo et al., 2018) randomized model parameters and training labels to test the sanity of attributions. However, this sanity check can lead to a functional mapping shift. Similar to the input perturbation-based techniques, sanity and sensitivity checks in columns 4-5 of Tab. 1 rely on input or model perturbations, making it challenging to simultaneously satisfy both input distribution invariance and functional mapping invariance.
Benchmark with Attribution Ground Truth. Other attribution benchmarks have made efforts to establish attribution ground truth for assessing attribution methods. Pointing Game (Zhang et al., 2018), DiFull & DiPart (Rao et al., 2022) utilize training annotations as attribution ground truth. However, training annotations only offer partial ground truth information for attributions. Khakzar et al. (Khakzar et al., 2022) calculated Null Feature without class information through model optimization as the ground truth for attributions. However, there remains a notable absence of evidence substantiating the relationship between the features generated by model optimization and their corresponding attributions. CLEVR-XAI (Arras et al., 2022) employs a visual question-answering framework to generate synthetic images to provide controlled evaluations of generated objects. However, the presence of shadows from these objects compromises the strict confinement of attribution to the objects’ physical areas. Lin et al. (Lin et al., 2021) employ neural Trojan for providing attribution ground truth, but their benchmark fails to guarantee that the model prediction change is solely attributable to Trojaned features. In addition to the partial guarantee of attribution verifiability, these attribution benchmarks evaluate the ratio between the defined ground truth region and the region with high attributions, compromising sensitivity in their evaluation metrics for attributions. From Tab. 1, it is evident that a full guarantee of attribution verifiability presents a significant challenge.
A.5 Fidelity Examination of Attribution Verifiability
To guarantee attribution verifiability, we carefully devise techniques for both backdoor and metric design to ensure that the prediction changes are fully attributed to the trigger features. During backdooring the model, we maintain low loss for both poisoned and clean samples while ensuring that trigger features substantially alter the prediction accuracy on poisoned samples, achieving up to 100% success as shown in Tabs. 3-5 of Supp. A.7.2.
[…]
A.6 Attribution Methods and Categories
In our experiments, nine attribution methods are benchmarked including GCAM (Selvaraju et al., 2017), FullGrad (Srinivas & Fleuret, 2019), Grad (Simonyan et al., 2014), GGCAM (Selvaraju et al., 2017), SG (Smilkov et al., 2017), IG (Sundararajan et al., 2017), IG-Uniform (Sturmfels et al., 2020), AGI (Pan et al., 2021), and LPI (Yang et al., 2023a). We categorize these methods into three groups; namely CAM-based, gradient-based, and integration-based methods, based on their distinct underlying explanation processes, as explained below.
CAM-based Methods. Zhou et al. (Zhou et al., 2016) proposed a class activation mapping (CAM) technique to localize the class-specific features of input samples. Assuming a feature map $A^{k}$ of the $k$-th convolutional layer before a softmax layer, CAM-based attribution methods $M^{c}_{CAM}$ (e.g., GCAM) calculate a weighted combination of activation maps $A^{k}$ for a class $c$ as
$M^{c}_{\text{CAM}}(x)=\Psi(ReLU(\sum_{k}w^{c}_{k}A^{k})),$ (9)
where $w^{c}_{k}=avg(\sum_{i}\sum_{j}\partial f_{c}(x)/\partial A^{k}_{i,j})$ indicates the activation weight by applying a global average pooling on the activation map $A^{k}$, and $\Psi$ indicates an interpolation operator to estimate an attribution map with the same scale as $x$ from a down-sampled activation map. In our experiments, we test two CAM-based attribution methods including GCAM and FullGrad. In contrast to CAM, GCAM avoid the use of additional structures enabling a generalized applications. FullGrad further integrates gradients of model biases to ensure the completeness axiom(Sundararajan et al., 2017). Since FullGrad retains a similar process of utilizing an interpolated activation map, we categorize it alongside GCAM as a CAM-based method. Although other CAM-based attribution methods also exist (Jiang et al., 2021; Muhammad & Yeasin, 2020), we test two representative ones in our experiments to cover this category.
Gradient-based Methods. Gradient-based attribution methods, such as Grad (Simonyan et al., 2014), employ a first-order Taylor expansion to approximate the non-linear model: $f_{c}(x)\approx w^{\intercal}*x+b$. Thus, the attribution of an input sample $x$ for a class $c$, can be computed using the gradients with respect to the input as
$M^{c}_{\text{Grad.}}(x)=\partial f_{c}(x)/\partial x.$ (10)
Integration-based Methods. Compared to gradient-based attribution methods, integration-based attribution methods integrate input gradients from a reference input $x^{\prime}$ to the feature $x$ along an integral path. Taking IG as an example, the attribution of an input feature $x_{i}$ for a class $c$ can be estimated as
$M^{c}_{i\ \text{Integr.}}(x,x^{\prime})=(x_{i}-x^{\prime}_{i})\times\int_{\alpha=0}^{1}\frac{\partial f_{c}(\tilde{x})}{\partial\tilde{x}_{i}}\bigg|_{\tilde{x}=x^{\prime}+\alpha(x-x^{\prime})}\mathrm{d}\alpha,\vskip 5.69054pt$ (11)
where $\alpha$ indicates a linear integration path from the reference input $x^{\prime}$ to the input $x$, and $x^{\prime}$ is typically set as a zero vector in IG. Different integration-based attribution methods (IG-Uni, IG-SG (Smilkov et al., 2017), AGI, and LPI) are proposed to redefine the reference and the integral path. As these methods share the approach of estimating integrals to offer explanations, we categorize these methods as integration-based attribution methods. Due to their distinct theoretical guarantees in using different references (Sundararajan et al., 2017; Yang et al., 2023a), we benchmark four integration-based attribution methods including IG, IG-Uni, AGI and LPI.
Other Methods. A variety of attribution methods also have been proposed for providing explanations through feature perturbation or removal (e.g., LIME (Ribeiro et al., 2016), SHAP (Lundberg & Lee, 2017), occlusion (Zeiler & Fergus, 2014), and mask (Fong & Vedaldi, 2017)). However, these perturbation-based methods are revealed to exhibit unreliabilities (Shrikumar et al., 2017; Sundararajan et al., 2017), as well as entailing a significant computational cost (Ancona et al., 2018). Therefore, we exclusively evaluate the propagation-based attribution methods. On the other hand, a few attribution methods, e.g., LRP (Binder et al., 2016) and DeepLift (Shrikumar et al., 2017), fail to satisfy basic axioms like completeness and implementation invariance. Hence, these methods are not included in our evaluation.
A.7 Experimental Setup
A.7.1 Trigger Patterns and Poisoned Samples
Related Works of Neural Trojans. Backdoors alter a model’s predictions by making it sensitive to certain trigger patterns. This framework is relevant because we leverage backdoors in our BackX. Gu et al. (Gu et al., 2019) introduced BadNet which mislabels and stamps triggers to poison the training samples. Blended attack (Chen et al., 2019), ISSBA (Li et al., 2021) and WaNet (Nguyen & Tran, 2021) are proposed to generate more stealthy sample-specific triggers, achieving invisible poisoning. Adap-Blend (Qi et al., 2022) further enhances the stealthiness of attack in the latent space. Due to the clear attribution of backdoor triggers in the models trained to achieve it, we leverage neural backdoors to circumvent the lack of ground truth in attribution evaluation. Given the wide applications of attribution methods in backdoor defense (Huang et al., 2019; Chou et al., 2020), we provide guidance for defense against backdoor attacks using attributions as part of this study.
Neural Trojan Setup for BackX. In our BackX benchmark, we incorporate both visible and invisible trigger patterns to provide a comprehensive evaluation, which allows us to thoroughly assess the capability of attribution methods in detecting both fixed visible patterns and input-specific invisible triggers. Figure 9 provides examples of the trigger patterns utilized in our experiments and the corresponding poisoned samples. In the Blend attack (Chen et al., 2019), we utilize fixed visible trigger patterns, consistent with Qi et al. (Qi et al., 2022). In addition, we employ a pre-trained encoder network to generate input-specific invisible trigger patterns in the ISSBA attack (Li et al., 2021). The trigger patterns used in these attacks are depicted in the middle column of Figure 9. Given a set of input samples, as shown in the first column of Figure 9, the resulting poisoned samples are presented in the last column of Figure 9.